OpenStudy (anonymous):

Which value of y is NOT part of the range of this function? f(x) = (9/x – 1) + 6

OpenStudy (johnweldon1993):

are there choices?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

I believe the range is \[(-\inf,6)U(6,\inf)\] due to the fact that if y were to = 6 you would get a false value.

OpenStudy (zzr0ck3r):

y = 9/(x-1) + 6 so (y-6)(x-1)=9 x=9/(y-6)1 thus y=6 is not defined

OpenStudy (anonymous):

Thank you both

OpenStudy (anonymous):

np

OpenStudy (zzr0ck3r):

bah that should say x=9/(y-6)+1

OpenStudy (zzr0ck3r):

to find range find the inverse(assuming there is one) then find its domain

OpenStudy (zzr0ck3r):

the domain of the inverse*